OpenStudy (anonymous):

Give the equation of a line that has a slope of 4 and passes through the point (3, –2). A. y = 4x – 14 B. y = –2x + 11 C. y = 3x – 14 D. y = –2x – 9

OpenStudy (anonymous):

if you plug in the slope and coordinates into \[y - y_{1}=m(x-x_{1})\] and solve for y you will have your y =mx +b equation :) i.e\[y +2 = 4(x-3)\] hope this was a nuff to get you started if not ask away :)

OpenStudy (anonymous):

or u can use y=mx+c m=slope

OpenStudy (anonymous):

yep thats the easy way ^ thanks for saying a alternate method XD

OpenStudy (anonymous):

so with y=mx+b it is -2 = 4(3) + b ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so -14 = b. what do I do next?

OpenStudy (anonymous):

Is A correctttt??

OpenStudy (anonymous):

yep..congrats..salu

OpenStudy (anonymous):

thanks a lot:)